OpenStudy (anonymous):

Nitrogen dioxide, a major air pollutant, can be produced by the combustion of nitrogen oxide as shown. 2NO + O2 -----> 2NO2 In a plant, 1,500 kg of nitrogen oxide is consumed per day to produce 1,500 kg of nitrogen dioxide per day. What is the percent yield? 21.7% 32.6% 43.5% 65.2%

OpenStudy (somy):

hmmm % yield= (actual/theoretical) * 100

OpenStudy (anonymous):

what is the actual and what is the theoretical though

OpenStudy (somy):

ok i think actual is the one the gave for NO2 =1500kg

OpenStudy (anonymous):

whats theoritcal

OpenStudy (somy):

mole of limiting reagent is compared with product mole

OpenStudy (somy):

by ratio

OpenStudy (somy):

mole = mass/ Mr mole = 1500 000/ 60 =25 000 mole (i converted kg to g)

OpenStudy (somy):

thats mole of NO

OpenStudy (somy):

ration between NO and O is 2:1 if 2---- 25000mole 1- x mole

OpenStudy (somy):

so x= 25000/2= 12 500mole for O2

OpenStudy (somy):

as u see O2 mole is less then of N so that means its limiting thus we use mole of O2 as the main mole to compare with the product

OpenStudy (somy):

ratio between O2 and NO2 is 1:2 1-- 12 500 mole 2 --- x x= 12 500* 2 = 25 000 mole is produced mass= mole * Mr mass = 25000* 92= 2300000 g /1000= 2 300 kg is your theoretical yield

OpenStudy (somy):

(1500/2300) *100= 65.2%

OpenStudy (somy):

so D im so sorry i guess u ran out of time?